For a given interest. Simple interest varies jointly as principle and time. IF $5000 is left in an account for 8 years and earns interest of $1600 how much interest would be earned if you deposited $10,000 for 7 years?
Answer by Theo(13342)

i believe the formula you are referring to is:
I = P * R * T
your first formula would be:
1600 = 5000 * 8 * R
you can use this formula to solve for R which will get you R = 1600 / (5000 * 8) which results in R = .04.
that's your annual interest rate.
4% per year for 8 year * 5000 principal gets you 1600 in interest.
now that you know R, you can solve for I in the second formula.
that formula becomes:
I = 10,000 * 7 * .04 which makes I = 2800.
the general formula would be:
I = P * R * T
In this formula, R is the constant of variation.
you used the first formula to solve for R and then you replaced R in the second formula to solve for I in that formula.
the first formula was:
1600 = 5000 * R * 8
the second formula, after solving for R in the first formula, was:
I = 10,000 * .04 * 7
solving for R in the first formula allowed you to solve for I in the second formula.